OpenStudy (anonymous):

How do you factor this expression as a single trigonometric function? cos x - sin²x - 1

zepdrix (zepdrix):

\[\Large\rm \cos x - \color{red}{\sin^2x} - 1\]You can do some fun stuff with this middle term, getting it in terms of cosine.

zepdrix (zepdrix):

Remember your Pythagorean Identity involving sine and cosine?

OpenStudy (anonymous):

I know sin²u + cos²u = 1. So sin²u = 1 - cos²u. Is that where you're going with this?

zepdrix (zepdrix):

Mmm yes good. Let's write our middle term in brackets like this, \[\Large\rm \cos x - (\color{red}{\sin^2x}) - 1\]It will make it easier to substitute in this way.

zepdrix (zepdrix):

\[\Large\rm \cos x - (\color{red}{1-\cos^2x}) - 1\]So that identity brings us here, yes?

OpenStudy (anonymous):

Alright, I get that. I'm not sure where to go from there.

zepdrix (zepdrix):

The reason I put brackets is because we need to distribute the negative to each term.\[\Large\rm \cos x -1+\cos^2x - 1\]

zepdrix (zepdrix):

Combining like-terms,\[\Large\rm \cos^2x+\cos x-2\]

zepdrix (zepdrix):

If you make a substitution something in your brain might click :)\[\Large\rm u=\cos x\]Giving us,\[\Large\rm u^2+u-2\]Any ideas? ^^ Stare at it for just a sec!!

OpenStudy (anonymous):

Can we factor it? (u - 1)(u +2)

zepdrix (zepdrix):

Mmm yes very good \c:/ From there just plug your cosx back in for u.

OpenStudy (anonymous):

so (cosx - 1)(cosx + 2) is the final answer?

zepdrix (zepdrix):

`factor this expression as a single trigonometric function` We have our factors, and it's in terms of a single trig function (cosine) Yes! Good job :)

OpenStudy (anonymous):

Thank you so much!